OpenStudy (anonymous):

-4(2x-9)+5(3x+2) can you please show me how to do this?

OpenStudy (waleed_imtiaz):

open the brackets..... -8x+36+15x+10=0 7x+46=0 x=-46/7

OpenStudy (anonymous):

can you please show me how to simplify it?

OpenStudy (waleed_imtiaz):

Yea.... sure...... Just as above I have solved it.... Apply PEMDAS rule because it involves paranthesis.... First open the brackets..... then add and subtract..... What's ur confusion ? Can U do this now ?

OpenStudy (anonymous):

I just don't know how you got all that... like i seen how you got -8x and 36 and 15x but how did u get 10? I am so confused!! Sorry.

OpenStudy (anonymous):

then 7x where did that come from?

OpenStudy (anonymous):

46 i know to added 36+10 to get 46.

OpenStudy (waleed_imtiaz):

oh...okay...... Just multiply -4 with (2x-9) and 5 with (3x+2)..... now U dont know from where 10 came..... multiply 5 with 3x and 5 with 2 U will get 15x+10.... Rite ? and the first bracket... -4 *2x and -4*-9...... after multiplying U will get .....-8x+36+15x+10=0 now just add 'x' terms...... U know 15x-8x=7x.... U got it ? And 36+10=46....... so to solve for x.... take 7 to the denominator of 46........ answer......

OpenStudy (anonymous):

oh okay. I think i got it. sorry and thanks. I have another one can you help me with it?

OpenStudy (waleed_imtiaz):

no problem.... Mm.... okay Just let me know whats that ...... :)

OpenStudy (anonymous):

5[2-3(6x-1)] this is exactly how it is in my book.

OpenStudy (waleed_imtiaz):

First of all ,we have to solve the parenthesis.... Okay....... So,,, Just keep everything as it is written..... and solve parenthesis.... =5[2-18x+3] =5[-18x+5] inside the bracket it is DMAS... SImple it is..... =-90x+25..... Now If on the other side it is equal to zero..... then u can find x.....U have any problem ?

OpenStudy (anonymous):

5*2-3*6x-1 is this what i am suppose to do first and then solve

OpenStudy (anonymous):

okay so how am i suppose to know when to do leave the parentheses and solve first or leave them like in this problem i needed to keep my parentheses?

OpenStudy (waleed_imtiaz):

@edmondss390 U just have to solve for parenthesis first.... See above I have multiplied and opened the parenthesis..... Then I do add and subtract in the square brackets and then in the end, multiply 5 with the remaining

OpenStudy (anonymous):

ok. thanks so much. I am sorry if i frustrated you. hope you have a wonderful day. thanks so much! :)

OpenStudy (waleed_imtiaz):

whatever wil happen, U just have to solve first the parenthesis, then braces and then square brackets......... and in each braket use DMAS rule.....its not that complicated, I am just explaining to much extent.....

OpenStudy (waleed_imtiaz):

no.,never..... I like mathematics... :) hehe.... thanks U.....